# A Global Minimum Tax for Large Firms Only:

### v2

@@ Abstract
Keywords: multinational firms; tax avoidance; global minimum tax; profit shifting; tax competition
@@ 1 Introduction
An important limitation of the GMT is, however, that it applies only to “large” multinationals, defined as multinational groups that have had no less than 750 million EUR in total annual revenues in at least two out of the last four years. [^3] Since 2016, multinational groups above this threshold size are also required to file individual reports on their activities, profits, and taxes paid in each of the countries in which they have a presence (country-by-country reporting). Figure 1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of all MNEs, which are responsible for about 90% of all MNE profits, have consolidated accounts above this threshold and are therefore covered by the GMT. [^4] This corresponds to the estimate of the OECD, which arrives at a coverage rate of 90% of profits, using the Orbis database and other data sources (OECD, 2020a, , # 505, p. 233). Nevertheless, a substantial amount of MNE profits exceeding 300 billion EUR remain outside the scope of the GMT. Moreover, the MNE sample in Orbis is not comprehensive and the database is known for oversampling large firms (Bajgar et al.,, 2020). Therefore, a coverage rate of the GMT of 90% of all MNE profits can be considered as an upper bound of the true profit share covered by the GMT.
@@ 2.1 Setup
We consider a tax competition model between a non-haven country, indexed by $n$, and a set of $H$ symmetric tax-haven countries, indexed by $h$. Tax competition arises between each of the tax havens and the non-haven. There are a large number of heterogeneous, multinational enterprises (MNEs) owned by individuals in the non-haven, each having a headquarters in the non-haven and an affiliate in each of the tax havens. The headquarters in the non-haven makes exogenous profits in this country, $\pi$, which are distributed in the range $[\underline{\pi},\infty)$, with a cumulative distribution function $F(\pi)$. Hence, all real activity occurs in the non-haven, and this is where the “true” profits accrue in its entirety. MNEs can shift a share $\theta\in[0,1]$ of their non-haven headquarters’ profits to the set of tax haven affiliates, in order to maximize their post-tax profits. [^8] This is the simplest possible setting for profit shifting. See e.g., Johannesen, (2022) or Krautheim and Schmidt-Eisenlohr, (2011) for a similar assumption. In particular, this setting avoids the possible non-existence of Nash equilibria that arise in models where the location of heterogeneous firms is endogenous (Baldwin and Okubo,, 2009; Davies and Eckel,, 2010; Haufler and Stähler,, 2013).
@@
The non-haven country maximizes a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of the heterogeneous multinationals, which are all owned by residents of the non-haven country. Since tax havens do not have any ownership in the firms’ profits, the welfare of each tax havens corresponds to their tax revenues.
We consider a three-stage game. In the first stage, the non-haven and haven countries simultaneously choose whether to commit to a single tax rate for all MNEs, or not. To comply with the GMT agreement, this single tax rate has to be equal to, or higher than, the GMT rate. [^10] This is related to the literature on endogenous timing, or leadership, in tax competition, where countries commit in a pre-play stage to set taxes early or late. See Kempf and Rota-Graziosi, (2010). In the second stage, countries that have committed to a single tax rate in the first stage set a single tax rate, whereas countries that have not committed choose different tax rates for large and small MNEs (where the tax rate on large MNEs has to be at least $t_{M}$). In the final stage, MNEs engage in profit shifting, given the tax rates in the non-haven and the symmetric tax havens.
In terms of tax choices, we make two assumptions. Firstly, countries can split their tax rate only if they are bound by the GMT. This reflects the fact that preferential tax regimes for specific tax base are effectively not allowed under the OECD’s action plan to fight base erosion (OECD,, 2013, Ch.3). Secondly, a set of tax havens decide whether to set a single GMT rate or split collectively. This assumption is just to simplify discussions and does not change the qualitative nature of tax-competition equilibrium or the quantitative results, which we will see below. We discuss an alternative decentralized decision process of tax havens in Appendix A.5. We solve this three-stage game by backward induction. The equilibrium concept we rely on is the sub-game perfect Nash equilibrium, which we call the tax-competition equilibrium.
@@ 2.3 Governments’ tax setting choices
Solving the first-order conditions of the two countries yields best responses: [^16] The second-order condition of both countries’ optimal tax problems are trivially satisfied.
@@ 2.4 Tax regimes
Table 1 shows that this sequence yields three potential choices for each of the two sets of countries. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In all these regimes there is a single tax in each haven country, which is at the same level or even above the GMT rate of the non-haven. Hence tax revenues in each haven country would be non positive. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven country for small MNEs. Hence only five possible tax regimes remain.
@@
In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates. To simplify our discussion, we treat the non-haven country as if it maximized tax revenues only. This is possible because the qualitative nature of our results does not depend on the marginal value of public funds, $\lambda$ [^18] Again, this differs from the results of Johannesen, (2022). The reason is that the ratio of marginal valuations for public vs. private consumption is constant in our analysis, and greater than one. All formal proofs are relegated to the Appendix.
@@
In Regime 1, where $t_{M}$ rises above the haven’s unconstrained tax rate $t_{h}^{0}$, the haven is forced to set $t_{M}$ for large MNEs. By contrast, the non-haven’s unconstrained rate is still above the GMT. As the non-haven is not bound by the GMT, it will continue to choose a single tax rate above the GMT for all MNEs. For the tax havens, the choice is therefore whether to apply the GMT also to small MNEs, or use a split tax schedule. By committing to a single GMT rate for all MNEs, the havens can induce the non-haven to set a higher uniform tax rate than in the unconstrained equilibrium. [^19] This is seen from the non-haven’s best response function (5), which is increasing in $t_{h}$. Havens therefore trade off the gains from inducing a higher uniform tax rate in the non-haven against the revenue losses from not undercutting the non-haven’s tax rate even more for small MNEs. In Regime 1, the first effect dominates and the havens find it profitable to commit to the GMT for all MNEs. In Regime 1, therefore, all countries choose higher uniform tax rates than they choose in the unconstrained equilibrium without a GMT.
@@
In Regime 2, where $t_{M}$ rises above the regime-switching rate $t_{M}^{1}$, the non-haven is still not bound by the GMT, and will continue to set a single tax rate above the GMT for all MNEs. For the havens, however, the trade-off described in Regime 1 above changes. As the GMT is now higher, there is less to gain from setting a uniform GMT rate for all MNEs, because the tax differential and thus shifted profits are reduced. [^20] This is seen from the tax rates in Regime 2 of Proposition 1: the tax differential, $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$, decreases in $t_{M}$. At the same time, given the higher GMT rate, there is more to gain for the havens from undercutting the non-haven’s tax rate for small MNEs. In Regime 2, therefore, the havens will find it optimal to split their tax rate and set $t_{M}$ only for the large MNEs, but choose a lower tax rate for small MNEs. [^21] In practice, this split is done by choosing the lower tax rate for small MNEs as the general corporate tax rate, but topping up this tax rate for large MNEs. See footnote 5 in the introduction. Johannesen, (2022) has shown formally that it is always rational for tax havens to use the QDMTT up to the total tax level of the GMT, rather than leaving the top-up taxation to the home countries of the MNEs. Note, finally, that the upper boundary of Regime 2, $t_{M}^{2}$, is rising in the GMT coverage rate, $\phi$. This is because the non-haven country faces less competition from the havens (i.e., higher tax rates $t_{h}$) for large firms. If large MNEs make up a large part of the corporate tax base, the non-haven country will therefore be able to set a uniform tax rate above $t_{M}$ for a wider range of GMT rates.
@@
Finally, in Regime 4, where $t_{M}>t_{M}^{3}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set the GMT rate for all firms. [^22] This is seen from the tax rates in Regime 3 of Proposition 1: the tax differential, $t_{n}^{3}-t_{h}^{3}=t_{M}/2$, increases in $t_{M}$. To avoid large tax base losses from small MNEs, the non-haven therefore abandons its commitment and also undercuts the GMT for small MNEs. Again, this argument is analogous to the one for the havens to end their commitment in Regime 2. In the Regime 4 equilibrium, both countries therefore split their tax rates and set their unconstrained tax rates for small MNEs, while adhering to the GMT for large MNEs.
@@ 3 Welfare effects of the GMT
Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of the introduction of the GMT and of two gradual reforms: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$. As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, including revenue maximization ($\lambda\rightarrow\infty$) as a special case. Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
@@ 3.1 Introduction of the GMT
In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0. As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs. In this case, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.242424If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits ($\phi\in(0,3/4)$, the introduction of the GMT rate will instead benefits the havens in Regime 3. See Appendix A.2 for details. Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes. These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting.
@@ 3.2 Gradual reforms of the GMT
Next we examine the welfare effects of the other reform option: a marginal increase in the GMT coverage rate $\phi$, while keeping the GMT rate $t_{M}$ fixed. As shown in (8), this is achieved by reducing the threshold profit level $\pi_{M}$ above which the GMT rate must be charged. The results are summarized in:
@@ 4 Quantitative implications
The most important result is that the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 2. In this regime, the havens find it profitable to split their tax rate and to set a lower tax rate for small MNEs as compared to large MNEs (Proposition 1(ii)). This corresponds to the observations mentioned in the Introduction that several tax havens have announced to maintain their low statutory tax rate while topping up the taxation of large MNEs to comply with the GMT, known as the QDMTT (see footnote 5). The GMT introduction reduces the total amount of profits shifted, thereby increasing the non-haven’s welfare while reducing the havens’ welfare (tax revenues).
@@
One may argue that the MVPF is indeed higher than our benchmark value of $\lambda=1.5$, as evidenced by the very substantial efforts, discussed in the introduction, that non-haven countries take to increase their corporate tax revenues. Panels (b) and (c) of Table 3 therefore present the results of two robustness checks with high values of the MVPF: $\lambda=3$ and $\lambda=10$. The latter value, given in Panel (c) approaches a Leviathan government objective of tax revenue maximization. With these increased values of $\lambda$, the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 1, in which both countries employ a uniform tax rate on all MNEs. With the higher valuation of tax revenues in the non-haven country, the tax havens will find it profitable to commit to a uniform GMT rate for all MNEs, which in turn leads the non-haven to choose a higher uniform tax rate above the GMT, as compared to the benchmark in Panel (a). The GMT rate of $15\%$ therefore acts as a coordination device between the non-haven and the havens, benefitting both countries. Although observations at the time of writing somewhat favor the GMT resulting in split tax rates in the havens, corresponding to Regime 2, we leave open the question of which regime will eventually emerge.
@@
Finally, Table 4 provides a quantitative assessment of two gradual reforms that are likely to follow the newly introduced GMT rate of $15\%$ with incomplete coverage. The first reform is to increase the GMT rate from $15\%$ to $18\%$ while maintaining the coverage rate at $\phi=90\%$ (Panel (a)). The second is to broaden the GMT coverage from $90\%$ to $100\%$ while keeping the GMT rate at $15\%$ (Panel (b)). In both scenarios, we assume our benchmark MVPF of $\lambda=1.5$ and thus, the equilibrium remains in Regime 2 after the reforms. The results indicate that the non-haven gains more from the $18\%$ GMT rate ($+1.0\%$ in welfare; $+7.0\%$ in revenues) than from the full GMT coverage ($+0.2\%$ in welfare; $+1.3\%$ in revenues). The greater gains for the non-haven from the $18\%$ GMT rate are driven by a much larger reductions in shifted profits to the havens ($-37.9\%$), as compared to the extension of GMT coverage ($-7.0\%$). Hence the increase in the GMT rate creates a much sharper conflict of interest with the tax havens, as compared to the extension of GMT coverage. In this sense, covering more MNEs under the GMT may be a politically more feasible option than increasing the GMT rate. However, even a full coverage of the GMT will likely bring only moderate gains to the non-haven and to the world, while entailing higher administrative and compliance costs.

### v3

@@ Abstract
Keywords: multinational firms; tax avoidance; global minimum tax; profit shifting
@@ 1 Introduction
An important limitation of the GMT is, however, that it applies only to “large” multinationals, defined as multinational groups that have had no less than 750 million EUR in total annual revenues in at least two out of the last four years. [^3] Since 2016, multinational groups above this threshold size are also required to file individual reports on their activities, profits, and taxes paid in each of the countries in which they have a presence (country-by-country reporting). Fig. 1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of all MNEs, which are responsible for about 90% of all MNE profits, have consolidated accounts above this threshold and are therefore covered by the GMT. [^4] This corresponds to the estimate of the OECD, which arrives at a coverage rate of 90% of profits, using the Orbis database and other data sources (OECD, 2020a, , # 505, p. 233). Nevertheless, a substantial amount of MNE profits exceeding 300 billion EUR remain outside the scope of the GMT. Moreover, the MNE sample in Orbis is not comprehensive and the database is known for oversampling large firms (Bajgar et al.,, 2020). Therefore, a coverage rate of the GMT of 90% of all MNE profits can be considered as an upper bound of the true profit share covered by the GMT.
@@ 2.1 Setup
We consider a tax competition model between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. Tax competition arises between each of the tax havens and the non-haven. There are a large number of heterogeneous, multinational enterprises (MNEs) owned by individuals in the non-haven, each having a headquarters in the non-haven and an affiliate in each of the tax havens. The headquarters in the non-haven makes exogenous profits in this country, $\pi$, which are distributed in the range $[\underline{\pi},\infty)$, with a cumulative distribution function $F(\pi)$. Hence, all real activity occurs in the non-haven, and this is where the “true” profits accrue in its entirety. MNEs can shift a share $\theta\in[0,1]$ of their non-haven headquarters’ profits to the set of tax haven affiliates, in order to maximize their post-tax profits. [^8] This is the simplest possible setting for profit shifting. See e.g., Johannesen, (2022) or Krautheim and Schmidt-Eisenlohr, (2011) for a similar assumption. In particular, this setting avoids the possible non-existence of Nash equilibria that arise in models where the location of heterogeneous firms is endogenous (Baldwin and Okubo,, 2009; Davies and Eckel,, 2010; Haufler and Stähler,, 2013).
@@
The non-haven country maximizes a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of the heterogeneous multinationals, which are all owned by residents of the non-haven country. Since tax havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
We consider a three-stage game. In the first stage, the non-haven and haven countries simultaneously choose whether to commit to a single tax rate for all MNEs, or not. [^10] This is related to the literature on endogenous timing, or leadership, in tax competition, where countries commit in a pre-play stage to set taxes early or late. See Kempf and Rota-Graziosi, (2010). To comply with the GMT agreement, this single tax rate has to be equal to, or higher than, the GMT rate. In our main model, we assume that the set of tax havens collectively decides on whether to set a single GMT rate, or split their tax rates. This assumption simplifies the algebra, but it does not change the qualitative nature of our results. [^11] We discuss the alternative of a decentralized first stage decision of the tax havens in Appendix A.2. In the second stage, all countries choose their tax rate(s) in a non-cooperative way. Countries that have committed to a single tax rate in the first stage set a single tax rate, whereas countries that have not committed choose different tax rates for large and small MNEs (where the tax rate on large MNEs has to be at least $t_{M}$). In the final stage, MNEs engage in profit shifting, given the tax rates in the non-haven and the symmetric tax havens. We solve this three-stage game by backward induction. The equilibrium concept we rely on is the sub-game perfect Nash equilibrium, which we call the tax-competition equilibrium.
@@ 2.3 Governments’ tax setting choices
Solving the first-order conditions of the two countries yields best responses: [^17] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
@@ 2.4 Tax regimes
Table 1 shows that this sequence yields three potential choices for each of the two sets of countries. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these potential regimes, if the non-haven sets a single GMT rate, the havens would choose a single tax that is at the same level or even higher. Hence tax revenues in each haven country would be zero. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven country for small MNEs. Alternatively, if the haven sets a single non-GMT or a single GMT tax rate, the non-haven will never want to split its tax rate, because the tax base elasticity of its MNEs is independent of profits. Hence only five possible tax regimes remain.
@@
In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates. Our discussion below will focus on tax revenue effects. This is strictly correct only for the tax havens. It is also possible for the non-haven, however, because the marginal value of public funds, $\lambda$, is constant in our analysis, and it always exceeds the valuation of private income. All formal proofs are relegated to the Appendix.
@@
In Regime 1, where $t_{M}$ rises above the haven’s unconstrained tax rate $t_{h}^{0}$, the haven is forced to set $t_{M}$ for large MNEs. By contrast, the non-haven’s unconstrained rate is still above the GMT. As the non-haven is not bound by the GMT, it will continue to choose a single tax rate above the GMT for all MNEs. For the tax havens, the choice is therefore whether to apply the GMT also to small MNEs, or use a split tax schedule. By committing to a single GMT rate for all MNEs, the havens can induce the non-haven to set a higher uniform tax rate than in the unconstrained equilibrium. This argument is strongest when tax havens collectively decide on whether to commit to a single tax rate, as we assume in our main model. [^19] This is seen from the non-haven’s best response function (5), which is increasing in $t_{h}$. However, as we show in Appendix A.2, the same qualitative argument also holds for decentralised decision-making of tax havens. Havens therefore trade off the gains from inducing a higher uniform tax rate in the non-haven against the revenue losses from not undercutting the non-haven’s tax rate even more for small MNEs. In Regime 1, the first effect dominates and the havens find it profitable to commit to the GMT for all MNEs. As a result, all countries choose higher uniform tax rates in Regime 1, as compared to the unconstrained equilibrium without a GMT.
@@
In Regime 2, where $t_{M}$ rises above the regime-switching rate $t_{M}^{1}$, the non-haven is still not bound by the GMT, and will continue to set a single tax rate above the GMT for all MNEs. For the havens, however, the trade-off described in Regime 1 above changes. As the GMT is now higher, there is less to gain from setting a uniform GMT rate for all MNEs, because the tax differential and thus shifted profits are reduced. [^20] This is seen from Regime 2 of Proposition 1, where the tax differential, $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases in $t_{M}$. At the same time, given the higher GMT rate, there is more to gain for the havens from undercutting the non-haven’s tax rate for small MNEs. In Regime 2, therefore, the havens will find it optimal to split their tax rate and set $t_{M}$ only for the large MNEs, but choose a lower tax rate for small MNEs. [^21] In practice, this split is done by choosing the lower tax rate for small MNEs as the general corporate tax rate, but topping up this tax rate for large MNEs. See footnote 5 in the introduction. Johannesen, (2022) has shown formally that it is always rational for tax havens to use the QDMTT up to the total tax level of the GMT, rather than leaving the top-up taxation to the home countries of the MNEs. Note, finally, that the upper boundary of Regime 2, $t_{M}^{2}$, is rising in the GMT coverage rate, $\phi$. This is because the non-haven country faces less competition from the havens (i.e., higher tax rates $t_{h}$) for large firms. If large MNEs make up a large part of the corporate tax base, the non-haven country will therefore be able to set a uniform tax rate above $t_{M}$ for a wider range of GMT rates.
@@
Finally, in Regime 4, where $t_{M}>t_{M}^{3}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set the GMT rate for all firms. [^22] This is seen from Regime 3 of Proposition 1, where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases in $t_{M}$. To avoid large tax base losses from small MNEs, the non-haven therefore abandons its commitment and also undercuts the GMT for small MNEs. Again, this argument is analogous to the one for the havens to end their commitment in Regime 2. In the Regime 4 equilibrium, both countries therefore split their tax rates and set their unconstrained tax rates for small MNEs, while adhering to the GMT for large MNEs.
@@ 3 Welfare effects of the GMT
Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of the introduction of the GMT and of two gradual reforms: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$. As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, and it includes revenue maximization ($\lambda\rightarrow\infty$) as a special case. Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
@@ 3.1 Introduction of the GMT
In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0. As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs. In this case, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.242424If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits so that $\phi\in(0,3/4)$, then the introduction of the GMT rate will instead benefit the havens in Regime 3. See Appendix A.3 for details. Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes. These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting.
@@ 3.2 Gradual reforms of the GMT
Next we examine the welfare effects of another likely reform option: a marginal increase in the GMT coverage rate $\phi$, while keeping the GMT rate $t_{M}$ fixed. As shown in (8), this is achieved by reducing the threshold profit level $\pi_{M}$ above which the GMT rate must be charged. The results are summarized in:
@@ 4 Quantitative implications
The most important result is that the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 2. In this regime, the havens find it profitable to split their tax rate and to set a lower tax rate for small MNEs as compared to large MNEs (Proposition 1(ii)). [^25] This result also holds if we assume that the first stage decision of whether to split tax rates or not is made by each tax haven individually (see Appendix A.2). In this case the range of GMT rates that support Regime 1 will be smaller than in our benchmark model, while the range of $t_{M}$ that supports an equilibrium in Regime 2 rises accordingly. This corresponds to the observations mentioned in the Introduction that several tax havens have announced to maintain their low statutory tax rate while topping up the taxation of large MNEs to comply with the GMT, known as the QDMTT (see footnote 5). The GMT introduction reduces the total amount of profits shifted, thereby increasing the non-haven’s welfare while reducing the havens’ welfare (tax revenues).
@@
One may argue that the MVPF is indeed higher than our benchmark value of $\lambda=1.5$, as evidenced by the very substantial efforts, discussed in the introduction, that non-haven countries take to increase their corporate tax revenues. Panels (b) and (c) of Table 3 therefore present the results of two robustness checks with high values of the MVPF: $\lambda=3$ and $\lambda=10$. The latter value, given in Panel (c) approaches a Leviathan government objective of tax revenue maximization. With these increased values of $\lambda$, the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 1, in which both countries employ a uniform tax rate on all MNEs. With the higher valuation of tax revenues in the non-haven country, the tax havens will find it profitable to commit to a uniform GMT rate for all MNEs, which in turn leads the non-haven to choose a higher uniform tax rate above the GMT, as compared to the benchmark in Panel (a). The GMT rate of $15\%$ therefore acts as a coordination device between the non-haven and the havens, benefitting both countries. [^27] These results depend on the assumption that tax havens make their first stage choice collectively. With decentralized first stage decisions of tax havens (cf. Appendix A.2), the upper boundary for a Regime 1 equilibrium, $t_{M}=t_{M}^{a}$, is calculated to be $9.37\%$, $9.42\%$, and $9.45\%$ for respective values of the marginal valuation of public funds, $\lambda=1.5$, $3$, and $10$. This implies that the introduction of a $15\%$ GMT rate leads to a Regime 2 equilibrium regardless of the level of $\lambda$. For $\lambda=1.5$, the quantitative effects under decentralized decision making of havens are the same as those in Panel (a) of Table 3. For $\lambda=3$ and $\lambda=10$, the effects differ from those in Panels (b) and (c) of Table 3, however, and yield a conflict of interests between the non-haven and the havens that is analogous to the one in Panel (a). Although observations at the time of writing somewhat favor the GMT resulting in split tax rates in the havens, corresponding to Regime 2, we leave open the question of which regime will eventually emerge.
@@
Finally, Table 4 provides a quantitative assessment of two gradual reforms that are likely to follow the newly introduced GMT rate of $15\%$ with incomplete coverage. The first reform is to increase the GMT rate from $15\%$ to $18\%$ while maintaining the coverage rate at $\phi=90\%$ (Panel (a)). The second is to broaden the GMT coverage from $90\%$ to $100\%$ while keeping the GMT rate at $15\%$ (Panel (b)). In both scenarios, we assume our benchmark MVPF of $\lambda=1.5$ and thus, the equilibrium remains in Regime 2 after the reforms. The results indicate that the non-haven gains more from the $18\%$ GMT rate ($+1.0\%$ in welfare; $+7.0\%$ in revenues) than from the full GMT coverage ($+0.2\%$ in welfare; $+1.3\%$ in revenues). The greater gains for the non-haven from the $18\%$ GMT rate are driven by the stronger reduction in profit shifting to the havens ($-37.9\%$), as compared to a full GMT coverage ($-7.0\%$). Hence the increase in the GMT rate creates a much sharper conflict of interest with the tax havens, as compared to the extension of GMT coverage. In this sense, covering more MNEs under the GMT may be a politically more feasible option than increasing the GMT rate. However, even a full coverage of the GMT will likely bring only moderate gains to the non-haven and to the world, while entailing higher administrative and compliance costs.